# Diffusion Models for Robotic Manipulation: A Survey

### v1

@@ 4.1.1 Imitation Learning
Although these methods are designed to enhance generalizability, achieving adaptability in highly dynamic environments and unfamiliar scenarios may require the integration of continuous and lifelong learning. This is a widely unexplored field in the context of DMs, with only very few works (Huang et al., 2024a, ; Di Palo et al.,, 2024) exploring this topic. Moreover, these methods are still limited in their applications. (Di Palo et al.,, 2024) are utilizing a lifelong buffer to accelerate the training of new policies for new tasks. In contrast, (Mendez-Mendez et al.,, 2023) continually updates their policy. However, they only conduct training and experiments in simulation. Additionally, their method requires precise feature descriptions of all involved objects and is limited to predefined abstract skills. Moreover, for the continual update, all past data is replayed, which is not only computationally inefficient but also does not prevent catastrophic forgetting.
@@
Addressing the disadvantage of long inference times with DMs, a novel approach employing flow matching has been introduced to generate actions for flexible and dynamic tasks, offering a robust alternative to traditional diffusion mechanisms (Black et al., 2024a, ). Compared to DP, which relies on an iterative denoising process, FM produces high-frequency action chunks for dexterous control by directly learning a continuous flow between states, thus avoiding the need for iterative refinement. Furthermore, unlike DP, which optimizes based on a denoising objective by progressively removing noise at each step, FM introduces a dual-loss strategy to enhance the model’s generalization capabilities. Specifically, tokens corresponding to continuous outputs are supervised using a flow matching loss, while tokens corresponding to discrete outputs are supervised using a cross-entropy loss.
VLAs offer access to models trained on huge amounts of data and with strong computational power, leading to strong generalization capabilities. To mitigate some of their shortcomings, such as imprecise actions, specialized policies can be used for refinement. To not restrict the generalizability of the VLA, DMs offer a great possibility, as they are able to capture complex multi-model distributions and process high-dimensional visual inputs. However, both VLAs and DMs have a relatively slow inference speed. Thus, especially in this combination with VLAs, increasing the sampling efficiency of DMs is important. One example was provided in the previous paragraph. But the topic of higher sampling speed with DMs is also discussed in more detail in section 2.2.1.
@@ 4.1.2 Offline Reinforcement Learning
To apply diffusion policies in the context of RL the reward term has to be integrated. Diffuser (Janner et al.,, 2022), one early work adapting diffusion to RL, uses classifier-based guidance, which is based on classifier guidance described in section 4.1.1. Let $\tau=\{(s_{0},a_{0}),\dots,(s_{T},a_{T})\}$ be a trajectory with one state-action pair per timestep in a planning horizon $\{0,\dots,T\}$. To incorporate the reward term during sampling, a regression model $R_{\phi}(\tau_{k})$ is trained to predicts the return, i.e., the cumulative future reward, over the trajectory $\tau_{k}$ at each denoising time step $k\in\{0,\dots,K\}$. This is incorporated into the sampling process, by adding the guidance term at each iteration of the reverse diffusion process (Janner et al.,, 2022):
@@
To mitigate these shortcomings, one approach, Decision Diffuser (Ajay et al.,, 2023), directly conditions the DM on the return of the trajectory using classifier-free guidance. This method outperforms Diffuser on a variety of tasks, such a a block-stacking task. However, both methods have not been evaluated on real-world tasks. Directly conditioning on the return, limits generalization capabilities. Different to Q-learning, where the value function is approximated, which generalizes across all future trajectories, here only the return of the current trajectory is considered. Sharing some similarity to on-policy methods, this limits generalization as the policy learns to follow trajectories from the demonstrations with high return values. Thus, this can also be interpreted as guided imitation learning.
A more common method (Wang et al., 2023a, ) integrates offline Q-learning with DMs. The loss function from eq. 5 is a behavior cloning loss, as the goal is to minimize error with respect to samples taken via the behavior policy. Wang et al., 2023a suggests including a critic in the training procedure, which they call Diffusion Q-learning (Diffusion-QL). In Diffusion-QL a Q-function is trained, by minimizing the Bellman-Operator using the double Q-learning trick. The actions for updating the Q-function are sampled from the DM. In turn a policy improvement step $\mathcal{L}_{c}=-\mathbb{E}_{\boldsymbol{s}\sim\mathcal{D},\boldsymbol{a}^{0}\sim\pi_{\theta}}\left[Q_{\phi}(\boldsymbol{s},\boldsymbol{a}^{0})\right]$ is included in the loss for updating the DM (Wang et al., 2023a, ):
@@
where $\mathcal{L}$ is the diffusion loss from eq. 5 and the parameter $\alpha$ regulates the influence of the critic. Several methods (Ada et al.,, 2024; Kim et al., 2024b, ; Venkatraman et al.,, 2023; Kang et al.,, 2023), build on Diffusion Q-learning. To increase the generalizability to out-of-distribution data, a common problem in offline RL (Levine et al.,, 2020), Ada et al., (2024), include a state-reconstruction loss, into the training of the DM.
@@ 4.2 Robotic grasp generation
Grasp generation with DMs can be categorized into several key approaches: From methodological perspective, one category focuses on explicit diffusion on 6-DoF grasp poses that lie on the $\mathbf{SE}(3)$ group, directly modeling spatial transformations to generate feasible grasps (Urain et al.,, 2023; Song et al., 2024b, ; Wu et al., 2024b, ; Weng et al.,, 2024; Singh et al.,, 2024; Lim et al.,, 2024). Another line of approaches involves implicit grasp diffusion within latent space, enhancing adaptability and versatility (Barad et al.,, 2024). A recent trend focuses on language-guided diffusion for task-oriented grasp generation, where natural language inputs shape the generation process (Nguyen et al., 2024a, ; Vuong et al.,, 2024; Nguyen et al., 2024b, ; Chang and Sun,, 2024). Other approaches emphasize affordance-driven diffusion, targeting specific functional goals, such as object pose diffusion for rearrangement (Liu et al., 2023b, ; Zhao et al.,, 2025), affordance-guided object reorientation (Mishra and Chen,, 2024), imitation learning (Wu et al., 2024a, ; Ma et al., 2024a, ) or multi-embodiment grasping (Freiberg et al.,, 2025). Apart from these categories, hand-object interaction (HOI) specifically prioritizes the synthesis of realistic, functional interactions by modeling the hand’s adaptive responses to various object shapes and affordances with dexterity (Ye et al.,, 2024; Wang et al., 2024c, ; Zhang et al., 2024d, ; Cao et al.,, 2024; Li et al., 2024a, ; Zhang et al.,, 2025; Lu et al.,, 2025; Zhang et al., 2024b, ). In addition to the diffusion on grasp generation or trajectory planning, DM as sim-to-real generator (Li et al., 2024d, ) or foundational feature extractor (Tsagkas et al.,, 2024) such as stable diffusion (Rombach et al., 2022a, ) may provide semantic information to enhance downstream grasp generation tasks. table 2 summarizes the aforementioned categories. Notably, we include the applications of diffusion in HOI, imitation learning for pre-grasp, and tasks related to image generation in the graph, which will not be further discussed in the rest of this survey due to their relevance to the field of computer vision. While readers are still encouraged to refer to the relevant literature according to our illustration (table 2: HOI Synthesis).
@@ 4.2.1 Diffusion as $\mathbf{SE}(3)$ grasp pose generation
To tackle this, $\mathbf{SE}(3)$-Diff (Urain et al.,, 2023) introduced a smooth cost function to learn the grasp quality via the energy-based model (EBM), where the score matching for EBM is applied on the Lie group to bridge the gap between diffusion processes on the vector space $\mathbb{R}^{6}$ and the $\mathbf{SE}(3)$. In contrast, (Song et al., 2024b, ) condition the 6-Dof grasp poses on the grasp locations $\mathbf{t}$ and corresponding volumetric features for grasp generation in clutter following GIGA framework (Jiang et al.,, 2021), without explicit consideration on the $\mathbf{SE}(3)$ constraint. Moreover, one advantage of the EBM model in $\mathbf{SE}(3)$-Diff is the direct grasp quality evaluation and integration into the entire grasp motion planning and optimization. However, training EBM-based models demands extensive sampling and poses significant challenges for generalization. We noticed that recent studies in Flow Matching (FM) (Lipman et al.,, 2023) is employed by EquiGraspFlow (Lim et al.,, 2024) and Grasp Diffusion Network (Carvalho et al.,, 2024), which usecontinuous normalizing flows (CNFs) as ODE solvers to learnangular ($\mathbf{SO}(3)$) and linear ($\mathbf{R}(3)$) velocities for denoising. This preserves the $\mathbf{SE}(3)$-equivariance conditioned on the input point cloud given the time schedule. In contrast to $\mathbf{SE}(3)$-Diff, which relies on additional supervision in the form of signed distance functions, they achieve competitive performance without requiring this auxiliary module, leading to more efficient training. In general, although CNFs-based approaches exhibits promising performance on grasp generation for a single object, more studies on the generalizability to highly-occluded environments (Freiberg et al.,, 2025; Shi et al.,, 2024) are expected in the future works.
@@ 5 Experiments and Benchmarks
In this section, we focus on the evaluation of the various DMs for robotic manipulation.
@@
Various benchmarks are used to evaluate the methods. Common benchmarks are CALVIN (Mees et al.,, 2022), RLBench (James et al.,, 2020), RelayKitchen (Gupta et al.,, 2020), and Meta-World (Yu et al.,, 2020). Primarily in RL, the benchmark D4RL Kitchen (Fu et al.,, 2020) is used. One method (Ren et al.,, 2024) uses FurnitureBench (Heo et al.,, ) for real-world manipulation tasks. Adroit (Rajeswaran et al.,, 2017) is a common benchmark for dexterous manipulation, LIBERO (Liu et al., 2023a, ) for lifelong learning, and LapGym (Maria Scheikl et al.,, 2023) for medical tasks.
@@
Many methods are only being evaluated against baselines, which are not based on DMs themselves. However, there are some common DM-based baselines. For methods operating in $\mathbf{SE}(3)$-space (Chen et al.,, 2024; Song et al., 2024b, ; Ryu et al.,, 2024), $\mathbf{SE}(3)$-Diffusion Policy (Urain et al.,, 2023), probably the first paper using DMs for grasp generation, is commonly used as baseline. For RL-based methods, the RL-based Diffuser (Janner et al.,, 2022), Diffusion-QL (Wang et al., 2023a, ), and Decision Diffuser (Ajay et al.,, 2023) are commonly used as baselines. It should be noted that in the original paper, Decision Diffuser (Ajay et al.,, 2023) is evaluated against Diffuser (Janner et al.,, 2022) and outperforms it on almost all tasks, particularly on the manipulation tasks, block stacking, and rearrangement. However, neither of these methods is evaluated on real-world tasks. Another common baseline is DP (Chi et al.,, 2023), as many methods are developed based on it. A common baseline for methods integrating 3D visual representations is 3D Diffusion Policy(Ze et al.,, 2024). 3D Diffusion Policy is evaluated against DP, and outperforms it on a huge variety of tasks in the benchmarks Adroit, MetaWorld, and Dexart with an average success rate of 74.4%, outperforming DP by 24.2%. It is also evaluated on four real-world manipulation tasks: rolling and pinching a dumpling, drilling, and pouring. With an average success rate of 85.0% it outperforms DP by 50%. 3D Diffusion Policy is greatly outperformed by 3D Diffuser Actor (Ke et al.,, 2024) on the CALVIN benchmark, especially for zero-shot long-horizon tasks. However, no comparison for real-world tasks is provided.
@@
The majority of methods are evaluated in simulation as well as in real-world experiments. For real-world experiments, most policies are directly trained on real-world data. However, some are trained exclusively in simulation and applied in the real world in a zero shot (Yu et al.,, 2023; Mishra et al.,, 2023; Ren et al.,, 2024; Liu et al., 2023b, ; Kapelyukh et al.,, 2024; Liu et al., 2023c, ), utilizing domain randomization, or real-world scene reconstruction in simulation. Few, predominately RL methods, are only evaluated in simulation (Yang et al.,, 2023; Power et al.,, 2023; Wang et al., 2023a, ; Janner et al.,, 2022; Pearce et al.,, 2022; Wang et al., 2023b, ; Mendez-Mendez et al.,, 2023; Kim et al., 2024b, ; Brehmer et al.,, 2023; Liang et al.,, 2023; Zhou et al., 2024a, ; Mishra and Chen,, 2024; Ajay et al.,, 2023; Ding and Jin,, 2023; Zhang et al., 2024a, ).

### v2

@@ 4.1.1 Imitation Learning
Although these methods are designed to enhance generalizability, achieving adaptability in highly dynamic environments and unfamiliar scenarios may require the integration of continuous and lifelong learning. This is a widely unexplored field in the context of DMs, with only very few works (Huang et al., 2024a, ; Di Palo et al.,, 2024) exploring this topic. Moreover, these methods are still limited in their applications. (Di Palo et al.,, 2024) are utilizing a lifelong buffer to accelerate the training of new policies for new tasks. In contrast, (Mendez-Mendez et al.,, 2023) continually updates its policy. However, they only conduct training and experiments in simulation. Additionally, their method requires precise feature descriptions of all involved objects and is limited to predefined abstract skills. Moreover, for the continual update, all past data is replayed, which is not only computationally inefficient but also does not prevent catastrophic forgetting.
@@
Addressing the disadvantage of long inference times with DMs, in some recent works instead, flow matching is used to generate actions from observations preprocessed by VLMs to solve flexible and dynamic tasks, offering a robust alternative to traditional diffusion mechanisms (Black et al., 2024a, ; Zhang and Gienger,, 2025). While Black et al., 2024a takes a skill-based approach, where the vision-language model is used to decide on actions, Zhang and Gienger, (2025) uses a vision-language model to generate waypoints. In both approaches, flow matching is used as the expert policy, generating precise trajectories.
VLAs offer access to models trained on huge amounts of data and with strong computational power, leading to strong generalization capabilities. To mitigate some of their shortcomings, such as imprecise actions, specialized policies can be used for refinement. To not restrict the generalizability of the VLA, DMs offer a great possibility, as they can capture complex multi-model distributions and process high-dimensional visual inputs. However, both VLAs and DMs have a relatively slow inference speed. Thus, especially in this combination with VLAs, increasing the sampling efficiency of DMs is important. One example was provided in the previous paragraph. But the topic of higher sampling speed with DMs is also discussed in more detail in Section 2.2.1.
@@ 4.1.2 Offline Reinforcement Learning
To apply diffusion policies in the context of RL the reward term has to be integrated. Diffuser (Janner et al.,, 2022), one early work adapting diffusion to RL, uses classifier-based guidance, which is based on classifier guidance described in Section 4.1.1. Let $\tau=\{(s_{0},a_{0}),\dots,(s_{T},a_{T})\}$ be a trajectory with one state-action pair per timestep in a planning horizon $\{0,\dots,T\}$. To incorporate the reward term during sampling, a regression model $R_{\phi}(\tau_{k})$ is trained to predict the return, i.e., the cumulative future reward, over the trajectory $\tau_{k}$ at each denoising time step $k\in\{0,\dots,K\}$. This is incorporated into the sampling process by adding the guidance term at each iteration of the reverse diffusion process (Janner et al.,, 2022):
@@
To mitigate these shortcomings, one approach, Decision Diffuser (Ajay et al.,, 2023), directly conditions the DM on the return of the trajectory using classifier-free guidance. This method outperforms Diffuser on a variety of tasks, such a block-stacking task. However, both methods have not been evaluated on real-world tasks. Directly conditioning on the return, limits generalization capabilities. Different to Q-learning, where the value function is approximated, which generalizes across all future trajectories, here only the return of the current trajectory is considered. Sharing some similarity to on-policy methods, this limits generalization as the policy learns to follow trajectories from the demonstrations with high return values. Thus, this can also be interpreted as guided imitation learning.
A more common method (Wang et al., 2023b, ) integrates offline Q-learning with DMs. The loss function from Eq. 5 is a behavior cloning loss, as the goal is to minimize error with respect to samples taken via the behavior policy. Wang et al., 2023b suggests including a critic in the training procedure, which they call Diffusion Q-learning (Diffusion-QL). In Diffusion-QL a Q-function is trained, by minimizing the Bellman-Operator using the double Q-learning trick. The actions for updating the Q-function are sampled from the DM. In turn a policy improvement step $\mathcal{L}_{c}=-\mathbb{E}_{\boldsymbol{s}\sim\mathcal{D},\boldsymbol{a}^{0}\sim\pi_{\theta}}\left[Q_{\phi}(\boldsymbol{s},\boldsymbol{a}^{0})\right]$ is included in the loss for updating the DM (Wang et al., 2023b, ):
@@
where $\mathcal{L}$ is the diffusion loss from Eq. 5 and the parameter $\alpha$ regulates the influence of the critic. Several methods (Ada et al.,, 2024; Kim et al., 2024b, ; Venkatraman et al.,, 2023; Kang et al.,, 2023), build on Diffusion Q-learning. To increase the generalizability to out-of-distribution data, a common problem in offline RL (Levine et al.,, 2020), Ada et al., (2024), include a state-reconstruction loss, into the training of the DM. An overview of the architectures of methods combining diffusion and reinforcement learning is provided in Table 3.
Reference Input Output Encoder Diffuser H/S
Janner et al., (2022) GTS RHC - U-Net
Ajay et al., (2023) GTS RHC - U-Net
Wang et al., 2023b GTS SiA - MLP
Wang et al., 2023c GTS RHC - DiT
Ding and Jin, (2023) GTS SiA - MLP
Mishra et al., (2023) GTS RHC - DiT
Kang et al., (2023) GTS RHC - MLP
Brehmer et al., (2023) GTS RHC - Eq. U-Net
Suh et al., (2023) GTS RHC - U-Net
Ha et al., (2023) RGBMV, Lan RHC ResNet, CLIP FiLM
Kim et al., 2024b GTS RHC - U-Net
Liang et al., (2023) GTS RHC - U-Net
Ada et al., (2024) GTS SiA - MLP
Ren et al., (2024) RGB/GTS SiA ViT/- U-Net/MLP
Huang et al., 2025b RGBSV SiA VQ-GAN VQ-Diffusion Gu et al., (2022)
Carvalho et al., (2023) GTS RHC - U-Net
Table 3: Technical details of trajectory diffusion using reinforcement learning. The references for the encoders are provided in Table 8.
In the following, the symbols and abbreviations are explained: H/S: Whether the method is hierarchical/skill-based (✓) or not (✗). Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), Eq.: Equivariant FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, Sia =single actions. A “-” indicates that no specialized encoder is required as ground truth state information is used.
@@ 4.2 Robotic grasp generation
Grasp generation with DMs can be categorized into several key approaches: From methodological perspective, one category focuses on explicit diffusion on 6-DoF grasp poses that lie on the $\mathbf{SE}(3)$ group, directly modeling spatial transformations to generate feasible grasps (Urain et al.,, 2023; Song et al., 2024b, ; Wu et al., 2024b, ; Weng et al.,, 2024; Singh et al.,, 2024; Lim et al.,, 2024). Another line of approaches involves implicit grasp diffusion within latent space, enhancing adaptability and versatility (Barad et al.,, 2024). A recent trend focuses on language-guided diffusion for task-oriented grasp generation, where natural language inputs shape the generation process (Nguyen et al., 2024a, ; Vuong et al.,, 2024; Nguyen et al., 2024b, ; Chang and Sun,, 2024). Other approaches emphasize affordance-driven diffusion, targeting specific functional goals, such as object pose diffusion for rearrangement (Liu et al., 2023b, ; Zhao et al.,, 2025), affordance-guided object reorientation (Mishra and Chen,, 2024), imitation learning (Wu et al., 2024a, ; Ma et al., 2024a, ) or multi-embodiment grasping (Freiberg et al.,, 2025). Apart from these categories, hand-object interaction (HOI) specifically prioritizes the synthesis of realistic, functional interactions by modeling the hand’s adaptive responses to various object shapes and affordances with dexterity (Ye et al.,, 2024; Wang et al., 2024c, ; Zhang et al., 2024d, ; Cao et al.,, 2024; Li et al., 2024a, ; Zhang et al.,, 2025; Lu et al.,, 2025; Zhang et al., 2024b, ). In addition to the diffusion on grasp generation or trajectory planning, DM as sim-to-real generator (Li et al., 2024d, ) or foundational feature extractor (Tsagkas et al.,, 2024) such as stable diffusion (Rombach et al., 2022a, ) may provide semantic information to enhance downstream grasp generation tasks. Table 4 summarizes the aforementioned categories. Notably, we include the applications of diffusion in HOI, imitation learning for pre-grasp, and tasks related to image generation in the graph, which will not be further discussed in the rest of this survey due to their relevance to the field of computer vision. While readers are still encouraged to refer to the relevant literature according to our illustration (Table 4: HOI Synthesis). More details on the architectures of the individual methods in grasp learning are provided in Table 5.
Reference Input Encoder Diffuser Benchmark
Urain et al., (2023) SDF Shape encoder FiLM Acronym
Barad et al., (2024) PCs PointNet++ FiLM Acronym
Song et al., 2024a TSDF OccNet FiLM VGN
Singh et al., (2024) PCs OccNet FiLM DA2
Lim et al., (2024) PCs VN-DGCNN FiLM Acronym
Freiberg et al., (2025) PCs + Gripper PCs Eq. U-Net Eq. FiLM Self generated
Carvalho et al., (2024) PCs PointNet++ DiT Acronym
Huang et al., 2025a PCs + Guidance VN-PointNet DiTs OakInk
Weng et al., (2024) PCs + Gripper PCs BPS DiTs DexGraspNet
Zhong and Allen-Blanchette, (2025) PCs + Gripper PCs Eq. Models Eq. DiTs MultiDex
Zhang et al., 2024h PCs PointNet++ DiTs MultiDex
Table 5: Technical details of grasp diffusion methodologies on $\mathbf{SE}(3)$ grasp synthesis. The references for the encoders are provided in Table 8. The references for the benchmarks are listed in Table 9.
In the following, the abbreviations used are explained: SDF: Signed Distance Function, TSDF: Truncated SDF, PCs: Point Clouds, FiLM: Convolutional Neural Network with Feature-wise Linear Modulation Perez et al., (2018), DiTs: Diffusion Transformers, Eq.: Equivariant, VN: Vector Neuron.
@@ 4.2.1 Diffusion as $\mathbf{SE}(3)$ grasp pose generation
To tackle this, $\mathbf{SE}(3)$-Diff (Urain et al.,, 2023) introduced a smooth cost function to learn the grasp quality via the energy-based model (EBM), where the score matching for EBM is applied on the Lie group to bridge the gap between diffusion processes on the vector space $\mathbb{R}^{6}$ and the $\mathbf{SE}(3)$. In contrast, (Song et al., 2024b, ) condition the 6-Dof grasp poses on the grasp locations $\mathbf{t}$ and corresponding volumetric features for grasp generation in clutter following GIGA framework (Jiang et al.,, 2021), without explicit consideration on the $\mathbf{SE}(3)$ constraint. Moreover, one advantage of the EBM model in $\mathbf{SE}(3)$-Diff is the direct grasp quality evaluation and integration into the entire grasp motion planning and optimization. However, training EBM-based models demands extensive sampling and poses significant challenges for generalization. We noticed that flow matching (Lipman et al.,, 2023) is employed in recent studies, such as EquiGraspFlow (Lim et al.,, 2024) and Grasp Diffusion Network (Carvalho et al.,, 2024), which use continuous normalizing flows (CNFs) as ODE solvers to learn angular ($\mathbf{SO}(3)$) and linear ($\mathbf{R}(3)$) velocities for denoising. This preserves the $\mathbf{SE}(3)$-equivariance conditioned on the input point cloud given the time schedule. In contrast to $\mathbf{SE}(3)$-Diff, which relies on additional supervision in the form of signed distance functions, they achieve competitive performance without requiring this auxiliary module, leading to more efficient training. In general, although CNF-based approaches exhibit promising performance on grasp generation for a single object, more studies on generalizability to highly occluded environments (Freiberg et al.,, 2025) and uncertainty quantification (Shi et al.,, 2024) are expected in future work.
@@ 5 Experiments and Benchmarks
In this section, we focus on the evaluation of the various DMs for robotic manipulation. Details on the employed benchmarks and baselines are listed in separate tables for imitation learning (Table 6), reinforcement learning (Table 7), and grasp learning (Table 5). Separately, the references for all applied benchmarks are listed in Table 9.
Reference Diffusion Baseline Simulation Real World
Benchmark #Demos Real #Demos
Diffusion Policy (DP) (Chi et al.,, 2023) FrankaKitchen, Robomimic, custom 566,
500,
ChainedDiffuser (Xian et al.,, 2023) RLBench 100 10 - 20
BESO (Reuss et al.,, 2023) DP, Diffusion-BC Relay Kitchen, CALVIN, custom 566,   /, 1000 -
Chen et al., 2023a CALVIN, FrankaKitchen, Ravens 200K,
566,
1000
Zhou et al., (2023) Diffuser∗1,
Decision Diffuser∗2
RLBench -
Diffusion-BC (Pearce et al.,, 2022) D4RLKitchen 566 -
Mendez-Mendez et al., (2023) BEHAVIOR - -
3D-DP (Ze et al.,, 2024) DP e.g. Adroit, MetaWorld, DexDeform 10 - 100 40
Ke et al., (2024) 3D-DP, ChainedDiffuser RLBench, CALVIN 24h 15
Liu et al., 2023c DP, SE(3)-DM custom / /
Power et al., (2023) custom / -
Ma et al., 2024b DP, Diffuser RLBench 100 20
Vosylius et al., (2024) DP RLBench 20 20
Zhang et al., 2024a Diffuser CALVIN / -
Reuss et al., (2024) CALVIN, LIBERO 24h, 50 4.5h
Scheikl et al., (2024) DP, BESO LapGym 90 - 200 90 - 200
Chen et al., (2024) DP, SE(3)-DM FrankaKitchen, Adroit 16k - 64k, 1.25k - 5k 60
Zhou et al., 2024a DP, BESO, Consistency Models∗3 Relay Kitchen, XArm Block Push, D3IL 566,
1k,
96 - 2k
Li et al., 2024c DP, 3D-DP Robomimic, custom 500,
100
100
Si et al., (2024) DP - 25 - 50
[…]
Bharadhwaj et al., 2024b EpicKitchens, RT1, BridgeData 400k ∗4 400
Wang et al., 2024b custom 50K trans∗5 50K trans∗5
Li et al., (2025) DP, 3D-DP RLBench 40 40
Table 6: Benchmarks of trajectory diffusion using imitation learning. For each benchmark, the numbers of demonstrations are listed in the same order. In the column “Diffusion Baselines” only those baselines, which are diffusion methods themselves, are listed. Methods not evaluated against a diffusion-based baseline, indicated by an (✗), are only evaluated against non-diffusion baselines or ablations of the method.
The references for the benchmarks are listed in Table 9. In the following, the symbols are explained: Methods by ∗1Janner et al., (2022), ∗2 Ajay et al., (2023), and ∗3 (Song et al.,, 2023). ∗4 The diffusion model is trained using uncurated video data. ∗5 As the number refers to the number of transitions, not demonstrations, this high number is expected. The column “Real” indicates whether methods are evaluated in the real world (✓), or not (✗). A “/” indicates that the information is not provided by the cited paper, while a “-” indicates that the information does not apply.
Reference Diffusion Baseline Simulation Real World
Benchmark #Demos Real #Demos
Diffuser (Janner et al.,, 2022) KUKA (custom) 10k -
Decision Diffuser (Ajay et al.,, 2023) D4RLKitchen, KUKA /, 10k -
Diffusion-QL (Wang et al., 2023b, ) D4RLKitchen 10000 trans∗1 -
Wang et al., 2023c Diffuser custom 8k -
HDMI (Li et al.,, 2023) - -
Ding and Jin, (2023) Diffusion-QL D4RLKitchen, Adroit / -
Mishra et al., (2023) Decision Diffuser STAP / /
Kang et al., (2023) Diffusion-QL Adroit, D4RL Kitchen / -
Brehmer et al., (2023) Diffuser KUKA / -
Suh et al., (2023) Diffuser - 100
Ha et al., (2023) Diffusion-QL - 50
Kim et al., 2024b Diffuser, Decision Diffuser, HDMI Fetch env / -
Liang et al., (2023) Diffuser, Decision Diffuser KUKA / -
Zhang et al., 2024a Diffuser CALVIN, CLEVR-Robot / -
Ada et al., (2024) Diffusion-QL / -
Ren et al., (2024) Diffusion-QL Robomimic,
D3IL,
FurnitureBench
100-300,
96,
50
∗2 50
Huang et al., 2025b Diffusion-QL MetaWorld, Adroit 20, 50 50
Carvalho et al., (2023) custom 25 -
Table 7: Benchmarks of trajectory diffusion using reinforcement learning. For each benchmark, the numbers of demonstrations are listed in the same order. In the column “Diffusion Baselines” only those baselines, which are diffusion methods themselves, are listed. Methods not evaluated against a diffusion-based baseline, indicated by an (✗), are only evaluated against non-diffusion baselines or ablations of the method. The references for the benchmarks are listed in Table 9. In the following, the symbols are explained: ∗1 As the number refers to the number of transitions, not demonstrations, this high number is expected. A (✓) in the column “Benchmark” indicates that the method is evaluated in simulation, but not with a robotic manipulation task, while a (✗) indicates that the method is not evaluated in simulation. The column “Real” indicates whether methods are evaluated in the real world (✓), or not (✗). A “/” indicates that the information is not provided by the cited paper, while a “-” indicates that the information does not apply.
@@
Various benchmarks are used to evaluate the methods. Common benchmarks are CALVIN (Mees et al., 2022b, ), RLBench (James et al.,, 2020), RelayKitchen (Gupta et al.,, 2020), and Meta-World (Yu et al.,, 2020). Primarily in RL, the benchmark D4RL Kitchen (Fu et al.,, 2020) is used. One method (Ren et al.,, 2024) uses FurnitureBench (Heo et al.,, ) for real-world manipulation tasks. Adroit (Rajeswaran et al.,, 2017) is a common benchmark for dexterous manipulation, LIBERO (Liu et al., 2023a, ) for lifelong learning, and LapGym (Maria Scheikl et al.,, 2023) for medical tasks.
@@
Many methods are only being evaluated against baselines, which are not based on DMs themselves. However, there are some common DM-based baselines. For methods operating in $\mathbf{SE}(3)$-space (Chen et al.,, 2024; Song et al., 2024b, ; Ryu et al.,, 2024), $\mathbf{SE}(3)$-Diffusion Policy (Urain et al.,, 2023), probably the first paper using DMs for grasp generation, is commonly used as baseline. For RL-based methods, the RL-based Diffuser (Janner et al.,, 2022), Diffusion-QL (Wang et al., 2023b, ), and Decision Diffuser (Ajay et al.,, 2023) are commonly used as baselines. It should be noted that in the original paper, Decision Diffuser (Ajay et al.,, 2023) is evaluated against Diffuser (Janner et al.,, 2022) and outperforms it on almost all tasks, particularly on the manipulation tasks, block stacking, and rearrangement. However, neither of these methods is evaluated on real-world tasks. Another common baseline is DP (Chi et al.,, 2023), as many methods are developed based on it. A common baseline for methods integrating 3D visual representations is 3D Diffusion Policy(Ze et al.,, 2024). 3D Diffusion Policy is evaluated against DP, and outperforms it on a huge variety of tasks in the benchmarks Adroit, MetaWorld, and Dexart with an average success rate of 74.4%, outperforming DP by 24.2%. It is also evaluated on four real-world manipulation tasks: rolling and pinching a dumpling, drilling, and pouring. With an average success rate of 85.0% it outperforms DP by 50%. 3D Diffusion Policy is greatly outperformed by 3D Diffuser Actor (Ke et al.,, 2024) on the CALVIN benchmark, especially for zero-shot long-horizon tasks. However, no comparison for real-world tasks is provided.
@@
The majority of methods are evaluated in simulation as well as in real-world experiments. For real-world experiments, most policies are directly trained on real-world data. However, some are trained exclusively in simulation and applied in the real world in a zero shot (Yu et al.,, 2023; Mishra et al.,, 2023; Ren et al.,, 2024; Liu et al., 2023b, ; Kapelyukh et al.,, 2024; Liu et al., 2023c, ), utilizing domain randomization, or real-world scene reconstruction in simulation. Few, predominately RL methods, are only evaluated in simulation (Yang et al.,, 2023; Power et al.,, 2023; Wang et al., 2023b, ; Janner et al.,, 2022; Pearce et al.,, 2022; Wang et al., 2023c, ; Mendez-Mendez et al.,, 2023; Kim et al., 2024b, ; Brehmer et al.,, 2023; Liang et al.,, 2023; Zhou et al., 2024a, ; Mishra and Chen,, 2024; Ajay et al.,, 2023; Ding and Jin,, 2023; Zhang et al., 2024a, ).